A car is traveling at a rate of 63 kilometer per hour. what is the car rate in Kilometer per min? how many K will the car travel in 10 minutes? Do not round.
Answer by davethejackal(28)

An hour is 60 minutes so you can travel 60x as far in an hour as you can in a minute ... so the amount you can travel in a minute is 1/60th that you can travel in an hour.
hence speed per minute = speed per hour / 60 = 63/60 = 21/20 = 1.05 km/min
In 10 minutes the car will travel 10x this distance = 1.05 x 10 km = 10.5km